A race car traveling northward on a straight, level track at constant speed travels 4,139 m in 180 seconds. The return trip over the same track is made in 165 seconds. a) What is the average speed of the car in m/s for the first half of the trip? b What is the average speed of the car over the entre trip?
Answer #1

a.)

Since, velocity(V) = (total distance travel(d))/(total taken time(t))

for first half,

given d1 = distance travel in first half trip = 4,139 m

t1 = time taken in first half trip = 180 sec.

So,

V1 = 4139/180

V1 = 22.99 m/sec.

(b.)

for entire trip,

Vavg = (d1 + d2)/(t1 + t2)

here, d2 = distance travel in return trip = d1 =4139 m

t2 = time taken in return trip = 165 sec.

So, Vavg = (4136 + 4136)/(180 + 165)

Vavg = 23.98 m/sec.

Please upvote.